Sorry to bother you with this but I'm stumped.
A city has an area of 4750 km^2. If the city were in the shape of a circle what would the diameter be?
Thank you in advance.

Answer by stanbon(48535)

A city has an area of 4750 km^2. If the city were in the shape of a circle what would the diameter be?
==============
Circle area = pi*r^2
pi*r^2 =4750
r^2 = 4750/pi = 1511.97
r = 38.88
diameter = 2r = 77.768 km
=============================
Cheers,
stan H.